OpenStudy (anonymous):

Find the slope of the line that passes through the pair of points. (2, –3), (–5, 1) Select one: a. 4/7 b. undefined c. 7/4 d. -7/4 e. -4/7

OpenStudy (anonymous):

slope = y2 - y1/x2 - x1 so the solve equals 1-(-3)/-5-2 =1+3/-5-2=4/-7=-4/7 the answer is E

OpenStudy (happy_to_help):

e will be correct because you need to use slope intercept form